OpenStudy (anonymous):

i hv been staring at these types of problems for a long time and im not really getting any of them rite...it would be nice if someone could help me.. thankx in advance..

OpenStudy (anonymous):

5 + 11 + 17 +....+ (6n-1) = n(3n + 2) 1) Show true for n=1 6(1) - 1 = (1)(3(1)+2) 6 - 1 = 1 (3 +2) 5 = 5 2) Show true for n=k 5 + 11 +17 + ...+ 6n - 1 = n (3n + 2) 5 + 11 + 17 + ...+(6k-1) = k(3(k)+2) 3)Show true for n = k+1 5+11 +17 +...+(6k - 1)+ (6(k+1)-1) = (k+1)(3(k+1)+2) k(3k+2) + 6(k) = (k+1)(3(k+1)+2) I dont know where to go from here....help..plz :(

OpenStudy (anonymous):

The directions are : use mathematical inductoin to prove that each statement is true for all natural numbers n

OpenStudy (anonymous):

That's not quite how you do induction. Your step 1 is right, but steps 2 and 3 are off. You need to prove that it's true for n = k+1, assuming that it is true for n = k. So, you want to set n = k and then prove it for n = k+1.

OpenStudy (anonymous):

Thats what i did..except i put "show" instead of "assume" :(

OpenStudy (anonymous):

can you help me..?

OpenStudy (anonymous):

Working on it, give me one minute.

OpenStudy (anonymous):

okay, no problem

OpenStudy (anonymous):

Oookay, I found the part that you messed up. You simplified (6(k+1)-1) to 6(k), but it should actually be 6k+5

OpenStudy (anonymous):

Other than that you're on the right track.

OpenStudy (anonymous):

\[ (6(k+1)-1) \to 6(k), \] simplify that!

OpenStudy (anonymous):

So @Shaq if we fix that part, we'll have \(k(3k+2)+(6k+5)=(k+1)(3(k+1)+2)\). Next step is to simplify that equation.

OpenStudy (anonymous):

omg thanku! i get it now; il ask if i need anyaditional help!

OpenStudy (anonymous):

Right on, any time :)

OpenStudy (anonymous):

so you always do the distribution before adding the additonal one?

OpenStudy (anonymous):

Remember order of operations, PEMDAS, Parentheses Exponents Multiplication/Division Addition/Subtraction. Parentheses always come first :)

OpenStudy (anonymous):

okay thanks! i think thats why i been getting these problems wrong. thanks a ton!

OpenStudy (anonymous):

Sure thing :)